OpenStudy (anonymous):

Please Help Me Question 11 (Multiple Choice Worth 1 points) (03.03) The equation below represents Function A and the graph represents Function B: Function A f(x) = -2x + 1 Function B graph of line going through ordered pairs negative 1, negative 5 and 2, 1 and 3, 3 Which equation best compares the slopes of the two functions? Slope of Function B = 2 × Slope of Function A. Slope of Function A = Slope of Function B Slope of Function A = 2 × Slope of Function B Slope of Function B = - Slope of Function A

OpenStudy (anonymous):

Find the slope for both function A and B and compare them.

OpenStudy (anonymous):

did u find the slope of the 2nd function??

OpenStudy (anonymous):

No im not good at these and I got to finish by 1:00

OpenStudy (anonymous):

u see, u have to put some effort...do u know how to find the slope of the 2nd line. anything in mind?

OpenStudy (anonymous):

make a function table. compare the x and y coordinates for function A.

OpenStudy (anonymous):

the slope is -2

OpenStudy (anonymous):

now compare that to the slope of function B. then you answer is the true statement.

OpenStudy (anonymous):

so its D

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Medal if helped?

OpenStudy (anonymous):

can you help me with more @GLKK

OpenStudy (anonymous):

@TylerSoSexy of course.

OpenStudy (anonymous):

Question 12 (Multiple Choice Worth 1 points) (03.03) The graph below shows four straight lines, P, Q, R, and S: graph of line P going through ordered pairs negative 2, negative 5 and 0, 3. Graph of line Q going through ordered pairs negative 1, negative 2 and 1, 6. Graph of line R going through ordered pairs negative 1, negative 4 and 1, 4. Graph of line S going through ordered pairs 0, negative 3 and 2, 5 Which line is represented by the function f(x) = 4x - 3? Line P Line Q Line R Line S Question 13 (Multiple Choice Worth 1 points) (03.01) Which of the following does not represent a function? graph of a negative parabola oriented about the y axis with y intercept at positive 6 graph of an absolute value function oriented about the y axis graph of an ellipse with x intercepts negative 8 and positive 8 graph of a negative parabola oriented about the y axis with y intercept at positive 6 Question 14 (Fill-In-The-Blank Worth 1 points) (03.05) The graph below plots a function f(x): graph of line segment going through ordered pairs 0,150 and 3, 0 0 If x represents time, the average rate of change of the function f(x) in the first three seconds is ___. Answer for Blank 1: Question 15 (Multiple Choice Worth 1 points) (03.02) Which of the following equations best represents the graph of a function f(m)? f(y) = m f(x) = m y = f(m) x = f(m)

OpenStudy (anonymous):

so for 12. make a function table for the function.

OpenStudy (anonymous):

the slope is 4 but I think its c

OpenStudy (anonymous):

it would look like this: x | f(x) = 4x - 3? | y 0 | 4*0-3 | 1 | 4*1-3 | 2 | 4*2-3 | 3 | 4*3-3 | The answers to the equation would be your y.

OpenStudy (anonymous):

@TylerSoSexy I would imagine C is correct. Fill out the function table and see what you get.

OpenStudy (anonymous):

its correct ii got it next one

OpenStudy (anonymous):

okay. question 13. it is C because functions graph coords not circles.

OpenStudy (anonymous):

question 14 I know you can do.

OpenStudy (anonymous):

is it three @GLKK

OpenStudy (anonymous):

@TylerSoSexy let me check.

OpenStudy (anonymous):

@TylerSoSexy I think it is but the question isn't too direct.

OpenStudy (anonymous):

okay it was next one I thank you for helping

OpenStudy (anonymous):

@TylerSoSexy you're welcome. 15 I would go with B.

OpenStudy (anonymous):

But I could be wrong again haha.

OpenStudy (anonymous):

B or D @TylerSoSexy

OpenStudy (anonymous):

can you help me with a couple more pleasssss

OpenStudy (anonymous):

Question 18 (Multiple Choice Worth 1 points) (03.06) Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3) Question 22 (Multiple Choice Worth 1 points) (03.07)Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4 Question 23 (Multiple Choice Worth 1 points) (03.07) Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -negative one halfx + 4. y = 2x + 8 y = 2x y = -2x - 8 y = -2x Question 24 (Multiple Choice Worth 1 points) (03.05) The graph of a function f(x) is shown below: graph of line going through ordered pairs -2, -1 and 3, 3. An open circle at -2, -1 and a black dot at 3, 3 What is the domain of f(x)? -2 < x ≤ 3 -2 ≤ x < 3 -1 < y ≤ 3 -1 ≤ y < 3 Question 25 (Multiple Choice Worth 1 points) (03.05) Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) + 4? The graph shifts 4 units up. The graph shifts 4 units down. The graph shifts 4 units left. The graph shifts 4 units right. Question 26 (Multiple Choice Worth 1 points) (03.05) The graph below shows the velocity f(t) of a runner during a certain time interval: graph of line going through ordered pairs 0, 4 and 4, 8. Graph of another line going through ordered pairs 4, 8 and 8, 0. Label on the x axis is time in second and label on the y axis is velocity in meter per second Which of the following describes the intercepts on the graph? The initial velocity of the runner was 4 m/s, and the runner stopped after 8 seconds. The initial velocity of the runner was 8 m/s, and the runner stopped after 4 seconds. The initial acceleration of the runner was 4 m/s2, and the runner stopped after 8 seconds. The initial acceleration of the runner was 8 m/s2, and the runner stopped after 4 seconds. Question 27 (Multiple Choice Worth 1 points) (03.05) The table below shows the distance d(t) in feet that an object travels in t seconds: t (second) d(t) (feet) 3 126 4 224 5 350 6 504 What is the average rate of change of d(t) between 3 seconds and 5 seconds, and what does it represent? 59.5 m/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 m/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 m/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds 59.5 m/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds Question 28 (Multiple Choice Worth 1 points) (03.03) Choose the equation that represents the graph below: Graph of a line passing through points 0 comma 4 and 3 comma 0 y = -negative three fourthsx + 4 y = -negative four thirdsx + 4 y = three fourthsx - 4 y = 4 thirdsx - 4 Question 29 (Multiple Choice Worth 1 points) (03.02) If f(x) = 2x2 + 2, find f(5). 102 52 22 -1.5 Question 30 (Multiple Choice Worth 1 points) (03.03) The location of an ant from a tree was 3 meters west (negative direction). The ant walked towards the tree at a constant rate and reached the tree in 2 minutes. If y represents the location of the ant from the tree, in meters, and x represents the time in minutes, which graph best represents this situation? graph of line going through ordered pairs negative 3, 0 and 0, 2 graph of line going through ordered pairs 0, negative 3 and 2, 0 graph of line going through ordered pairs 0, 2 and 3, 0 graph of line going through ordered pairs 0, negative 2 and negative 3, 0

OpenStudy (anonymous):

@GLKK @lxl-rhea-lxl

OpenStudy (anonymous):

@TylerSoSexy I have to go eat Lunch. I could try to help when I get back in an hour.

OpenStudy (anonymous):

@tylercollins

OpenStudy (anonymous):

@trotay97

OpenStudy (anonymous):

@crabbypattyking

OpenStudy (anonymous):

Your cheating. I'm in FLVS Algebra one and this is the practice exam.

OpenStudy (anonymous):

@bloofoffiction your cheating too lol but i am too sorta when i ask for help i dont ask for answers i ask for an explanation